A is 7 year older to B. 15 years ago B's age was three fourth of A's age. Find their ages in years now
Answer by stanbon(75887)

A is 7 year older to B. 15 years ago B's age was three fourth of A's age. Find their ages in years now
----
Equations:
A = B + 7
B-15 = (3/4)(A-15)
----
Substitute for "A" and solve for "B":
B-15 = (3/4)(B+7-15)
-----
4(B-15) = 3(B-8)
4B-60 = 3B-24
B = 36 yrs (B's age now)
---
A = B + 7 = 43 yrs (A's age now)
--------
Cheers,
Stan H.
--------